# Data-Centric Evolution in Autonomous Driving: A Comprehensive Survey of Big Data System, Data Mining, and Closed-Loop Technologies

### 23 Jan 2024 v1

@@ Abstract
The aspiration of the next generation’s autonomous driving (AD) technology relies on the dedicated integration and interaction among intelligent perception, prediction, planning, and low-level control. There has been a huge bottleneck regarding the upper bound of autonomous driving algorithm performance, a consensus from academia and industry believes that the key to surmount the bottleneck lies in data-centric autonomous driving technology. Recent advancement in AD simulation, closed-loop model training, and AD big data engine have gained some valuable experience. However, there is a lack of systematic knowledge and deep understanding regarding how to build efficient data-centric AD technology for AD algorithm self-evolution and better AD big data accumulation. To fill in the identified research gaps, this article will closely focus on reviewing the state-of-the-art data-driven autonomous driving technologies, with an emphasis on the comprehensive taxonomy of autonomous driving datasets characterized by milestone generations, key features, data acquisition settings, etc. Furthermore, we provide a systematic review of the existing benchmark closed-loop AD big data pipelines from the industrial frontier, including the procedure of closed-loop frameworks, key technologies, and empirical studies. Finally, the future directions, potential applications, limitations and concerns are discussed to arouse efforts from both academia and industry for promoting the further development of autonomous driving.
@@ 1 Background Introduction and Motivation
The advancement of deep learning technology combined with hardware supercomputing capacity and big data collection ability, has aroused the emergence of autonomous driving algorithms, which involve a range of task modules such as detection Wu et al. (2023b); Hou et al. (2023), tracking Zhang et al. (2023b); Li and Jin (2022), trajectory prediction Wang et al. (2022); Zhang et al. (2023a), occupancy forecasting Khurana et al. (2023); Zheng et al. (2023), motion planning Huang et al. (2023); Chen et al. (2023a), and low-level control Kim et al. (2022). To achieve the ultimate goal of safe, adaptive, and efficient autonomous driving, all of the algorithms and softwares must be built upon a highly automated closed-loop data driven pipeline Zhang et al. (2022a); Chen et al. (2023b) for self-evolution and upgrading. This is the pre-requisite of the predicted large-scale commercial application of autonomous vehicle (AV) technology in year 2025 Khan et al. (2022). Despite extensive research works have done to solely focus on designing AD algorithm itself using either open-source datasets or self-collected datasets, little efforts have been devoted to developing a systematic technology pipeline of data-centric autonomous driving, which in turn give negative impacts to the entire systems’ performance improvement, resulting in the existing bottleneck of AD algorithm performance upper-bound. The above mentioned concerns motivate us to carry out a comprehensive review regarding the state-of-the-art data-centric autonomous driving technologies, including AD big data system, AD data mining, and closed-loop AD technologies.
@@
Data is the new oil in autonomous driving world. There has been a severe Long-Tail Distribution problem in real-world autonomous driving (AD) applications Long et al. (2022). In order to fully address the long-tail distribution problem, it is estimated that we need to accumulate more than 1000 Billion driving mileage’s big data Zhou et al. (2023), which is far from being realized yet. Then, it comes to the big question: how to solve the long-tail distribution problem? There are three insights regarding to the answer. First, existing rule-based methodologies fail to address the problem, even in planning and decision-making tasks which are previously considered as their strength Lan et al. (2023). Second, closed-loop data-driven approach is indispensable for advanced AD algorithm development and deployment, with an emphasize on automatically alleviating long-tail distribution problem. Third, we should re-consider the way of collecting, storing, and utilizing the massive autonomous driving data. Data collection should across all types of essential sensors equipped on AVs, not only camera recorded driving videos. During the storage and utilization procedure, information privacy, anonymity, and security should be guaranteed.
@@
There are a few pioneer industrial practices in data-driven autonomous driving Farabet and Koumchatzky (2020); Elluswamy (2023); Gao (2023). Among them, Tesla is a representative, with its long-term developed Fleet Learning pipeline M (2022) and AutoPilot system. Whenever a corner case is detected by the AutoPilot, there will be a quick ”snapshot” record of the case, which comprised of all the major sensors’ sequential data during 1-minute time. After receiving the snapshot record, it will be analyzed by DL-based models and/or human experts to extract the features of the scenario, then the ”Shadow Mode” Silva et al. (2022) will be opened on a large number of vehicles to automatically find and record driving scenario data which are highly-similar in feature dimension with the previous corner case. The new added huge amount data will be in turn feed back to the AD algorithms for model training and validation, facilitating the algorithm’s upgrade and improvement.
@@
Meanwhile, the research community also closely focuses on summarizing the main technology routes and pipelines of autonomous driving. Ma et al. (2023); Li et al. (2023b) presented systematic reviews of AD perception algorithms, Huang et al. (2022) and Ma et al. (2022) presented a comprehensive survey of AD prediction/planning algorithms, respectively. In the perspective of data-driven autonomous driving, Li et al. (2023a); Liu et al. (2024) summarized the development of mainstream autonomous driving datasets. For closed-loop technology, Zhang et al. (2022b) investigated how it could be used for intelligent vehicle planning based on reinforcement learning (RL) methods. However, existing works only focus on one aspect of AD big data technology, none of them conducted a comprehensive review on the whole horizon of data-centric autonomous driving, including the dataset system development, data mining technologies, closed-loop technologies, and empirical studies of data-driven autonomous driving.
@@
  • Providing a systematic review of the state-of-the-art closed-loop data-driven autonomous driving pipelines and related key technologies based on deep learning and generative AI.
@@
The systematic organization of this paper is shown in Figure 1. We start by providing a comprehensive review regarding the development and key features of benchmark autonomous driving datasets in Section 2. Then, Section 3 introduces the state-of-the-art closed-loop data-driven autonomous driving system, including the pipelines, key technologies, and industrial empirical studies. Following that, the pros and cons of existing methodologies and pipelines are discussed in Section 4, with a guidance of future research direction. Finally, we conclude this paper in Section 5.
@@ 2 State-of-the-art Autonomous Driving Datasets: Classification and Development
The evolution of autonomous driving datasets mirrors the technological advancements and growing ambitions in the field. Early endeavors in the late 20th century, such as the MIT’s AVT Research Fridman et al. (2019) and UC Berkeley’s PATH Program Shladover et al. (1991), laid the groundwork with basic sensor data, but were limited by the technology level of the era. There has been a significant leap forward over the last two decades, fueled by advancements in sensor technology, computational power, and sophisticated machine learning algorithms. In 2014, the Society of Automotive Engineers (SAE) announced a systematic six-level (L0-L5) autonomous driving system to the public Ding et al. (2023), which has been widely recognized by autonomous driving R&D progress. Empowered by deep learning, computer vision-based methods have been dominant for intelligent perception. Deep reinforcement learning and its variants have provided crucial improvements in intelligent planning and decision-making. More recently, Large Language Models (LLMs) and Vision Language Models (VLMs) showcase their strong capability of scene understanding, driving behavior reasoning & prediction, intelligent decision making, which open up new possibilities for future development of autonomous driving.
@@ 2.1 Milestone Development of Autonomous Driving Datasets
Figure 2 shows the milestone development of the open-source autonomous driving datasets following the chronological order. According to notable developments, the mainstream datasets can be categorized into three identical generations, marked by significant leaps in dataset complexity, volume, scenario variety, and labeling granularity, propelling the field into new frontiers of technology sophistication. Specifically, the horizontal axis denotes the development timeline. The side header of each row includes dataset name, sensor modality, suitable tasks, places of data collection, and related challenges. To further compare datasets across different generations, we visualize the perception and prediction/planning dataset scales using bar chart with different color. The early phase, denoted as the 1st generation begin at 2012, was spearheaded by KITTI Geiger et al. (2012) and Cityscapes Cordts et al. (2016), which provided high-resolution images for perception tasks and were fundamental in benchmarking progress of visual algorithms. Advancing to the 2nd generation, datasets such as NuScenes Caesar et al. (2020), Waymo Sun et al. (2020), Argoverse 1 Chang et al. (2019) introduced a multi-sensor approach, integrating vehicle camera, High-Definition Maps (HD-Map), LiDAR, Radar, GPS, IMU, trajectory, surrounding agents’ data together, which was crucial for comprehensive driving environment modeling and decision-making processes.
@@
Moving towards recently, NuPlan Caesar et al. (2021), Argoverse 2 Wilson et al. (2021), Lyft L5 Houston et al. (2021) have significantly raised the impact bar, offering unprecedented data scales and fostering an ecosystem conducive to cutting-edge research. These datasets, characterized by their vast scale and multi-modal sensor integration, have been instrumental in developing algorithms for perception, prediction, and planning tasks, which paving the way for the state-of-the-art End2End or hybrid autonomous driving models. Coming to year 2024, we are now welcoming the 3rd generation autonomous driving datasets. Empowered by VLMs Wen et al. (2023); Cui et al. (2024), LLMs Qian et al. (2023); Wu et al. (2023a), and other Generative AI technologies, the 3rd generation datasets underscore the industry’s commitment to addressing the increasingly complex challenges of autonomous driving, such as data Long-Tail Distribution problem, Out-of-Distribution (OOD) detection, Corner Case Analysis, etc.
@@ 2.2 Dataset Acquisition, Settings, and Key Features
For prediction and planning tasks, we summarize the input/output components, sensor suite, scene length, and prediction length of the mainstream datasets in Table 2. As for motion prediction/forecasting task, the input components usually consists of ego-vehicle historical trajectory, surrounding agent historical trajectory, HD-Map, and traffic state information (i.e., traffic signal status, road ID, stop signs, etc). The target output is several most possible trajectories (such as the Top-5 or Top-10 trajectories) of ego-vehicle and/or surrounding agents in a short period of time. Motion forecasting task often adopts sliding time-window setting, where the whole scene is divided into several shorter time-windows. For instance, NuScenes adopts the past 2-second ground-truth data and HD-Map to predict the next 6-second, whereas Argoverse 2 takes the historical 5-second ground-truth and HD-Map to forecast the future 6-second trajectory. NuPlan, CARLA, and ApolloScape are among the most popular datasets for planning task. The input components consists of ego/surrounding vehicle historical trajectories, ego-vehicle motion states, and driving scenario representations. While NuPlan and ApolloScape are acquired in real-world, CARLA is a simulated dataset. CARLA contains road images captured during simulated driving sessions in different towns. Each road image i accompanied with a steering angle, which represents the required adjustment to keep the vehicle on normal driving. The prediction length of planning can be varied according to the requirements of different algorithms.
@@
Weather Time-of-day Driving Condition Camera LiDAR GPS&IMU CAN-bus HD-Map
KITTI 2012 1 1 3 FV 2D b-box &3D b-box
CityScapes 2016 1 1 3 FV 2D Seg
Mapillary 2016 5 3 3 SV 2D Seg
Apolloscape 2016 4 3 3 FV 3D b-box & 2D Seg
BDD 2018 5 2 4 FV Text
SemanticKITTI 2019 1 1 4 $_$ 3D Seg
DAWN 2019 6 3 3 FV 2D b-box
UNDD 2019 2 2 2 FV 2D Seg
@@
DriveLM Dataset. Autonomous driving systems are progressing, yet they still lack critical capabilities essential for robust operation in real-world, such as accurately handle unseen scenarios or unfamiliar objects, interact with human drivers, provide explanations in decision-making progress. As a pioneer of exploring the 3rd generation’s autonomous driving big data, DriveLM aims to tackle these challenges by incorporating VLMs,LLMs, and Graph Visual Question Answering (GVQA) technologies into the AD stack. DriveLM leverages these strengths to foster generalization in planning and to facilitate human-robot interaction through models that can understand and generate natural language descriptions. This compatibility with human communication is a key differentiator from current methods that rely on purely end-to-end differentiable architectures. DriveLM is particularly innovative in the design of thinking chain GVQA under the AD context. GVQA involves formulating Pi-3 reasoning—consisting of a series of question-answer pairs—to simulate the human reasoning process involved in driving. Aside from that, DirveLM also introduces dedicated metrics for assessing performance in a more logical and reasoned manner.

### 26 Jan 2024 v2

@@ Abstract
The aspiration of the next generation’s autonomous driving (AD) technology relies on the dedicated integration and interaction among intelligent perception, prediction, planning, and low-level control. There has been a huge bottleneck regarding the upper bound of autonomous driving algorithm performance, a consensus from academia and industry believes that the key to surmount the bottleneck lies in data-centric autonomous driving technology. Recent advancement in AD simulation, closed-loop model training, and AD big data engine have gained some valuable experience. However, there is a lack of systematic knowledge and deep understanding regarding how to build efficient data-centric AD technology for AD algorithm self-evolution and better AD big data accumulation. To fill in the identified research gaps, this article will closely focus on reviewing the state-of-the-art data-driven autonomous driving technologies, with an emphasis on the comprehensive taxonomy of autonomous driving datasets characterized by milestone generations, key features, data acquisition settings, etc. Furthermore, we provide a systematic review of the existing benchmark closed-loop AD big data pipelines from the industrial frontier, including the procedure of closed-loop frameworks, key technologies, and empirical studies. Finally, the future directions, potential applications, limitations and concerns are discussed to arouse efforts from both academia and industry for promoting the further development of autonomous driving. The project repository is available at: https://github.com/LincanLi98/Awesome-Data-Centric-Autonomous-Driving.
@@ 1 Background Introduction and Motivation
The advancement of deep learning technology combined with hardware supercomputing capacity and big data collection ability, has aroused the emergence of autonomous driving algorithms, which involve a range of task modules such as detection Wu et al. (2023b); Hou et al. (2023), tracking Zhang et al. (2023b); Li and Jin (2022), occupancy forecasting Khurana et al. (2023); Zheng et al. (2023), trajectory prediction Wang et al. (2022); Zhang et al. (2023a), motion planning Huang et al. (2023); Chen et al. (2023a), and low-level control Kim et al. (2022). To achieve the ultimate goal of safe, adaptive, and efficient autonomous driving, all of the algorithms and softwares must be built upon a highly automated closed-loop data driven pipeline Zhang et al. (2022a); Chen et al. (2023b) for self-evolution and upgrading. This is the pre-requisite of the predicted large-scale commercial application of autonomous vehicle (AV) technology in year 2025 Khan et al. (2022). Despite extensive research works have done to solely focus on designing AD algorithm itself using either open-source datasets or self-collected datasets, little efforts have been devoted to developing a systematic technology pipeline of data-centric autonomous driving, which in turn give negative impacts to the entire systems’ performance improvement, resulting in the existing bottleneck of AD algorithm performance upper-bound. The above mentioned concerns motivate us to carry out a comprehensive review regarding the state-of-the-art data-centric autonomous driving technologies, including AD big data system, AD data mining, and closed-loop AD technologies. By adopting this research, we anticipate that it will enhance the industrial R&D process and algorithm performance of autonomous driving, while also providing valuable insights to related academic studies in the same field.
@@
Data is the new oil in autonomous driving world. There has been a severe Long-Tail Distribution problem in real-world autonomous driving applications Long et al. (2022). In order to fully address the long-tail distribution problem, it is estimated that we need to accumulate more than 1000 Billion driving mileage’s big data Zhou et al. (2023), which is far from being realized yet. Therefore, the big question arises: how to solve the long-tail distribution problem? There are three insights regarding to the answer. First, when encounter an unfamiliar condition/corner case, AD algorithms become unreliable and may make wrong decisions. However, such instances are exceedingly rare in the database, and are challenging to be acquired from real-world. Under this circumstance, AD data mining and advanced data generation technologies hold promise for addressing the problem. Second, closed-loop data-driven approaches are indispensable for advanced AD algorithm development and deployment, with an emphasize on automatically alleviating long-tail distribution problem. Third, we should re-consider the way of collecting, storing, and utilizing the massive autonomous driving data. Data should be collected across all types of essential sensors equipped on AVs, not only camera recorded driving videos. During the storage and utilization procedure, information privacy, anonymity, and security should be guaranteed.
@@
There are a few pioneer industrial practices in data-driven autonomous driving Farabet and Koumchatzky (2020); Elluswamy (2023); Gao (2023). Among them, Tesla is a representative, with its long-term developed Fleet Learning pipeline M (2022) and AutoPilot system. Whenever a corner case is detected by the AutoPilot, there will be a quick ”snapshot” record of the case, which is comprised of all the major sensors’ data during 1-minute time. After receiving the snapshot record, it will be analyzed by DL-based models and/or human experts to extract the features of the scenario, then the ”Shadow Mode” Silva et al. (2022) will be opened on a large number of vehicles to automatically find and record driving scenario data which are highly-similar in feature dimension with the previous corner case. The newly added huge amount data will be feed back to the AD algorithms for model training and validation, facilitating the algorithm’s upgrade and improvement.
@@
Meanwhile, the research community also closely focuses on summarizing the main technology routes and pipelines of autonomous driving. Ma et al. (2023); Li et al. (2023b) presented systematic reviews of AD perception algorithms, Huang et al. (2022) and Ma et al. (2022) presented a comprehensive survey of AD prediction/planning algorithms, respectively. From the perspective of data-driven autonomous driving, Li et al. (2023a); Liu et al. (2024) summarized the development of mainstream autonomous driving datasets. For closed-loop technology, Zhang et al. (2022b) investigated how it could be used for intelligent vehicle planning based on reinforcement learning (RL) methods. However, existing works only focus on one aspect of AD big data technology; none of them conducted a comprehensive review on the whole horizon of data-centric autonomous driving, including the dataset system development, data mining technologies, closed-loop technologies, and empirical studies of data-driven autonomous driving.
@@
  • Providing a systematic review of the state-of-the-art closed-loop data-driven autonomous driving pipelines and related key technologies based on deep learning and generative AI models.
@@
The systematic organization of this paper is shown in Figure 1. We start by providing a comprehensive review regarding the development and key features of benchmark autonomous driving datasets in Section 2. Then, Section 3 introduces the state-of-the-art closed-loop data-driven autonomous driving system, including the pipelines, key technologies, and industrial empirical studies. Following that, the pros and cons of existing methodologies and pipelines are discussed in Section 4, with guidance on future research directions. Finally, we conclude this paper in Section 5.
@@ 2 State-of-the-art Autonomous Driving Datasets: Classification and Development
The evolution of autonomous driving datasets mirrors the technological advancements and growing ambitions in the field. Early endeavors in the late 20th century, such as the MIT’s AVT Research Fridman et al. (2019) and UC Berkeley’s PATH Program Shladover et al. (1991), laid the groundwork with basic sensor data, but were limited by the technology level of the era. There has been a significant leap forward over the last two decades, fueled by advancements in sensor technology, computational power, and sophisticated machine learning algorithms. In 2014, the Society of Automotive Engineers (SAE) announced a systematic six-level (L0-L5) autonomous driving system to the public Ding et al. (2023), which has been widely recognized by autonomous driving R&D progress. Empowered by deep learning, computer vision-based methods have dominated intelligent perception. Deep reinforcement learning and its variants have provided crucial improvements in intelligent planning and decision-making. More recently, Large Language Models (LLMs) and Vision Language Models (VLMs) showcase their strong capability of scene understanding, driving behavior reasoning & prediction, and intelligent decision making, which open up new possibilities for future development of autonomous driving.
@@ 2.1 Milestone Development of Autonomous Driving Datasets
Figure 2 illustrates the milestone development of the open-source autonomous driving datasets following the chronological order. Notable advancements have led to the categorization of mainstream datasets into three identical generations, marked by significant leaps in dataset complexity, volume, scenario variety, and labeling granularity, propelling the field into new frontiers of technological sophistication. Specifically, the horizontal axis denotes the development timeline. The side header of each row includes dataset name, sensor modality, suitable tasks, places of data collection, and related challenges. To further compare datasets across different generations, we visualize the perception and prediction/planning dataset scales using bar charts with different colors. The early phase, denoted as the 1st generation begin at 2012, was spearheaded by KITTI Geiger et al. (2012) and Cityscapes Cordts et al. (2016), which provided high-resolution images for perception tasks and were fundamental in benchmarking progress of visual algorithms. Advancing to the 2nd generation, datasets such as NuScenes Caesar et al. (2020), Waymo Sun et al. (2020), Argoverse 1 Chang et al. (2019) introduced a multi-sensor approach, integrating vehicle camera, High-Definition Maps (HD-Map), LiDAR, Radar, GPS, IMU, trajectory, surrounding agents’ data together, which was crucial for comprehensive driving environment modeling and decision-making processes.
@@
Moving towards recently, NuPlan Caesar et al. (2021), Argoverse 2 Wilson et al. (2021), and Lyft L5 Houston et al. (2021) have significantly raised the impact bar, offering unprecedented data scales and fostering an ecosystem conducive to cutting-edge research. These datasets, characterized by their vast scale and multi-modal sensor integration, have been instrumental in developing algorithms for perception, prediction, and planning tasks, paving the way for the state-of-the-art End2End or hybrid autonomous driving models. Coming to year 2024, we are now welcoming the 3rd generation of autonomous driving datasets. Empowered by VLMs Wen et al. (2023); Cui et al. (2024), LLMs Qian et al. (2023); Wu et al. (2023a), and other Generative AI technologies, the 3rd generation datasets underscore the industry’s commitment to addressing the increasingly complex challenges of autonomous driving, such as data Long-Tail Distribution problem, Out-of-Distribution (OOD) detection, Corner Case Analysis, etc.
@@ 2.2 Dataset Acquisition, Settings, and Key Features
For prediction and planning tasks, we summarize the input/output components, sensor suite, scene length, and prediction length of the mainstream datasets in Table 2. As for motion prediction/forecasting task, the input components usually consist of ego-vehicle historical trajectory, surrounding agent historical trajectory, HD-Map, and traffic state information (i.e., traffic signal status, road ID, stop signs, etc). The target output is several most possible trajectories (such as the Top-5 or Top-10 trajectories) of ego-vehicle and/or surrounding agents in a short period of time. Motion forecasting task often adopts sliding time-window settings, where the whole scene is divided into several shorter time-windows. For instance, NuScenes adopts the past 2-second ground-truth data and HD-Map to predict the next 6-second, whereas Argoverse 2 takes the historical 5-second ground-truth and HD-Map to forecast the future 6-second trajectory. NuPlan, CARLA, and ApolloScape are among the most popular datasets for planning task. The input components consist of ego/surrounding vehicle historical trajectories, ego-vehicle motion states, and driving scenario representations. While NuPlan and ApolloScape are acquired in real-world, CARLA is a simulated dataset. CARLA contains road images captured during simulated driving sessions in different towns. Each road image i accompanied with a steering angle, which represents the required adjustment to keep the vehicle on normal driving. The prediction length of planning can be varied according to the requirements of different algorithms.
@@
Weather Time-of-day
Drive
Condition
Region Camera LiDAR GPS&IMU CAN-bus HD-Map
KITTI[1] 2012 1 1 3 EU FV 2D b-box &3D b-box
CityScapes[2] 2016 1 1 3 EU FV 2D Seg
Mapillary[3] 2016 5 3 3 Global SV 2D Seg
Apolloscape[4] 2016 4 3 3 AS FV 3D b-box & 2D Seg
BDD[5] 2018 5 2 4 NA FV Text
SemanticKITTI[6] 2019 1 1 4 EU $_$ 3D Seg
DAWN[7] 2019 6 3 3 Unknown FV 2D b-box
UNDD[8] 2019 2 2 2 Unknown FV 2D Seg
@@
  • [1] Geiger et al. (2012), [2] Cordts et al. (2016), [3] Neuhold et al. (2017), [4] Huang et al. (2018), [5] Yu et al. (2020), [6] Behley et al. (2019), [7] Kenk and Hassaballah (2020), [8] Nag et al. (2019), [9] Caesar et al. (2020), [10] Chang et al. (2019), [11] Sun et al. (2020), [12] Liao et al. (2022), [13] Mao et al. (2021), [14] Caesar et al. (2021), [15] Wilson et al. (2021), [16] Sima et al. (2023)
@@
  • [1] Wilson et al. (2021), [2] Caesar et al. (2020), [3] Ettinger et al. (2021), [4] Huang et al. (2018), [5] Caesar et al. (2021), [6] Dosovitskiy et al. (2017)
@@
DriveLM Dataset. Autonomous driving systems are progressing, yet they still lack critical capabilities essential for robust operation in the real world, such as accurately handling unseen scenarios or unfamiliar objects, interacting with human drivers, providing explanations in decision-making progress. As a pioneer of exploring the 3rd generation’s AD big data, DriveLM aims to tackle these challenges by incorporating VLMs, LLMs, and Graph Visual Question Answering (GVQA) technologies into the AD stack. DriveLM leverages these strengths to foster generalization in planning and to facilitate human-robot interaction through models that can understand and generate natural language descriptions. This compatibility with human communication is a key differentiator from current methods that rely on purely end-to-end differentiable architectures. DriveLM is particularly innovative in the design of thinking chain GVQA under the AD context. GVQA involves formulating Pi-3 reasoning—consisting of a series of question-answer pairs—to simulate the human reasoning process involved in driving. Aside from that, DirveLM also introduces dedicated metrics for assessing performance in a more logical and reasoned manner.